EVA Report – January 10th

SOL: 09

Earth date: 10/01/2017

Person filling out the report: Gwendal Hénaff

Crew members involved in the EVA: Gwendal Hénaff, Patrick Gray, Pierrick Loyers

EVA leader: Gwendal Hénaff

Begin: 14:30 am

End: 17:00 am

Type of EVA: Driving and Hiking

 Purpose: 3D cartography measurements

UTM Coordinates:          Northing: 4252073 Easting: 520369 Zone/Sector: 12S

Northing: 4252066 Easting: 517460 Zone/Sector: 12S


Due to the bad weather conditions and following Shannon’s advice, we postponed the EVA to the afternoon. The objective was to reach Green Mars View but due to the risk of rain, we decided to stop closer, nearby Dinosaur Quarry. We get the necessary elevation measurements to scan the subsurface in the next days. Everything went fine. We went back to the hab 2 hours and 30 minutes after the beginning of the EVA. We noticed lot of mud on both Deimos and the ATV due to the wet areas we had to cross on the road. The sun was already low, and we decided to postpone the cleaning to tomorrow morning.